# Nonlinear Landau levels in the almost-bosonic anyon gas

### v1

@@ Abstract
We consider the quantitative description of a many-particle gas of interacting abelian anyons in the plane, confined in a trapping potential. If the anyons are modeled as bosons with a magnetic flux attachment, and if the total magnetic flux is small compared to the number of particles, then an average-field description becomes appropriate for the low-energy collective state of the gas. Namely, by means of a Hartree–Jastrow ansatz, we derive a two-parameter Chern–Simons–Schrödinger energy functional which extends the well-known Gross–Pitaevskii / nonlinear Schrödinger density functional theory to the magnetic (anyonic) self-interaction. One parameter determines the total number of self-generated magnetic flux units in the system, and the other the effective strength of spin-orbit self-interaction. This latter interaction can be either attractive/focusing or repulsive/defocusing, and depends both on the intrinsic spin-orbit interaction and the relative length scale of the flux profile of the anyons. Densities and energies of ground and excited states are studied analytically and numerically for a wide range of the parameters and align well with a sequence of exact nonlinear Landau levels describing Jackiw–Pi self-dual solitons. With increasing flux, counter-rotating vortices are formed, enhancing the stability of the gas against collapse. Apart from clarifying the relations between various different anyon models that have appeared in the literature, our analysis sheds considerable new light on the many-anyon spectral problem, and also exemplifies a novel supersymmetry-breaking phenomenon.
@@
$\int_{D(0,1)}\left[|\nabla f|^{2}+\alpha^{2}\frac{|\mathbf{x}|^{2}}{\max\{R,|\mathbf{x}|\}^{4}}|f|^{2}+\alpha g\frac{{\mathds{1}}_{B(0,R)}}{R^{2}}|f|^{2}\right],$
@@
where $P,Q$ are linearly independent and coprime complex polynomials with $n:=\max(\deg P,\deg Q)=\beta/2$. This complete set of regular minimizers includes all the self-dual soliton solutions found by Jackiw and Pi: initially, radial vortex rings with $P=z^{n}$ and $Q=1$ JacPi-90b , and eventually, the more general rationals $P/Q$ JacPi-90a . Thus, along the critical line $\gamma=-2\pi\beta$, $\beta\in\mathbb{R}$, the sequence of points $\beta=2n$ mark a phase transition from strictly positive (broken supersymmetry) to zero energy and increasing degree of supersymmetry, exhibited by NLL manifolds of real dimension $4n$; compare Figs. 2 and 3 (top).
@@
For a general estimate of the ground-state energy at $\gamma=2\pi\beta\kappa$, $\kappa\geq-1$, we have compared with the linear interpolation $E^{\mathrm{TF}}(2\pi c(1+\kappa)\beta)=E^{\mathrm{TF}}(2\pi\beta)\sqrt{c}\sqrt{1+\kappa}$ (dash-dotted curve), which provides a fit to numerics to within a few percent for $\beta\gtrsim 3$ and $0.9\lesssim\kappa\lesssim 2$. The estimate given by ChiSen-92 for $\kappa=1$ is ca 6% smaller. An absolute lower bound for the energy (assuming regularity and singlevaluedness) is provided by $\gamma_{*}(\beta)\to 2\pi\beta$ and thus differs by a factor $\sqrt{c}$, i.e. 9% (dotted curve; which for ideal anyons is still 41% higher than the ideal Fermi energy). We also conjecture that the real-valued minimizer at $\kappa=\pm 1$ has a linearly interpolating energy $|2\pm\beta|$. This, as well as other linear features of the $\kappa=\pm 1$ spectra, resembles the known linear/noninterpolating exact states of the $N$-anyon spectrum Chou-91a ; MurLawBhaDat-92 ; ChiSen-92 ; AtaGirLun-25 .
@@ Acknowledgments.
We are grateful to Romain Duboscq for the original Matlab code upon which much of our numerical work is based. A.A., A.E., F.G. and D.L. thank the Department of Mathematics of Politecnico di Milano for its kind hospitality during the spring of 2025 intensive period “Quantum Mathematics at Polimi”. D.L. also thanks the GSSI L’Aquila for its hospitality. All of us have benefited from discussions at the Institut Mittag-Leffler workshop “Anyons from small to large scales” in July 2025. Financial support from the Swedish Research Council (D.L., grant no. 2021-05328, “Mathematics of anyons and intermediate quantum statistics”) is gratefully acknowledged. A.A. thanks Kaj Nyström for the financial support for traveling to the relevant conferences. The author T.G. acknowledges the Gran Sasso Science Institute for the financial support making this work possible. A.E. gratefully thanks the Center for Interdisciplinary Mathematics at Uppsala University as well as C.F. Liljewalchs stipendiestiftelse for their financial support.
@@
Due to the determinantal property of the Wronskian and the symmetry of the amplitude $|P|^{2}+|Q|^{2}$, the corresponding NLL state $\phi_{P,Q}$ given by (8) is invariant under pointwise action of the group $\mathrm{SU}(2)\times\mathbb{R}^{+}$ on the space of polynomial pairs $(P,Q)$. It follows that we may without loss of generality assume that $P$ or $Q$ is monic, and, e.g., that $\deg Q<\deg P=n=\beta/2$. We may also simplify the problem (15) by adding any multiple of $Q$ to $P$. If we identify states that differ only by multiplication by a constant phase factor $e^{i\theta}$, we have a further $\mathrm{U}(1)$-symmetry, for a full $\mathrm{U}(2)\times\mathbb{R}^{+}$-symmetry in total. See AtaLunNgu-24 for further details.
@@
with $w_{1},w_{2}\in\mathbb{C}$. If $w_{1}=w_{2}$, this is a vortex ring of degree 2. For $d=3$, we have two particular solutions:
@@
For $d=4$, we have the particular solutions
@@
These particular solutions yield all NLL states for $n=3$ via (16).
@@
First we make some general observations. By expanding the kinetic-magnetic term $\lvert(\nabla+i\mathbf{A}_{\varrho})\phi\rvert^{2}$, the energy functional can be recast into the form
@@
which satisfies $\int_{\mathbb{R}^{2}}\varrho^{a}=1$. The corresponding magnetic potential is
@@
$\int_{\mathbb{R}^{2}}\lvert\mathbf{A}_{\varrho^{a}}\rvert^{2}\varrho^{a}=\vartheta(a),$
where
$\vartheta(a):=(2a-1)\left(2\psi(4a-1)-\psi(2a)-\psi(6a-2)\right).$ (46)
This can be shown by evaluating
$\int_{\mathbb{R}^{2}}\lvert\mathbf{A}_{\varrho^{a}}\rvert^{2}\varrho^{a}=2(2a-1)\int_{0}^{\infty}\frac{(1-(1+r^{2})^{-(2a-1)})^{2}}{r(1+r^{2})^{2a}}\,dr$
@@
$=\frac{1}{\sqrt{a-1}}\sqrt{\frac{4a(2a-1)}{(2a+1)}+\frac{1}{\pi}\frac{2a-1}{4a-1}\gamma+2\vartheta(a)\beta^{2}}.$ (48)
For fixed $(\beta,\gamma)$, this expression can be numerically optimized over $a>1$. This has been done using Mathematica, and the resulting curve for $\gamma=2\pi\beta\kappa$ is the thick yellow line in the plots.
We note that the Gaussian can be re-obtained by e.g. setting $a=2/b$, $\lambda=\sqrt{b/2}$, and taking the limit $b\to 0$. Hence the state $\phi_{a}$ can (for optimal $a>1$) be expected to have a lower energy than the Gaussian state.
@@
Here we tabulate approximate energies of (2) computed in Matlab and optimized using a steepest descent algorithm adopted from CorDubLunRou-19 . As initial states we have started from the states generated in the work CorDubLunRou-19 with $\gamma=0$, and adjusted the parameters $(\beta,\gamma)$ to nearby values in order to generate approximate ground states with different vorticities (the number of vortices listed is approximate in non-symmetric cases). In the stable regime away from criticality, convergence to the desired precision is typically achieved after 50-200 iterations of the algorithm. Due to geometrically and topologically induced metastability, the energy converges into a local minimum, and the global minimum may only be found by comparing a set of reasonable vorticities. Some chosen initial states exhibit sufficient metastability that they can produce estimates for excited states well above the ground levels, by stopping the algorithm after a limited number of iterations during which a stable phase is shown. Thus, all numerically computed energies are only valid as approximate upper bounds to the true ground-state energy, but may also yield indications of higher resonances. At criticality $\kappa=-1$ we have only been able to obtain sufficiently stable and reliable numerics for radially symmetric states. We have also noted some observations, such as approximate linearity, comparison to analytical energies [^1] Note that the constant $c$ was denoted by $e(1,1)/(2\pi)$ in the earlier work CorDubLunRou-19 ., or cases where we suspect that the state is far from converged. The resulting energies are shown in the plots using solid markers (colors and shapes indicate numbers of visible vortices; see legend in Fig. 4 (bottom)).

### v2

@@ Abstract
We consider the quantitative description of a many-particle gas of interacting abelian anyons in the plane, confined in a trapping potential. If the anyons are modeled as bosons with a magnetic flux attachment, and if the total magnetic flux is small compared to the number of particles, then an average-field description becomes appropriate for the low-energy collective state of the gas. Namely, by means of a Hartree–Jastrow ansatz, we derive a two-parameter Chern–Simons–Schrödinger energy functional which extends the well-known Gross–Pitaevskii / nonlinear Schrödinger density functional theory to the magnetic (anyonic) self-interaction. One parameter determines the total number of self-generated magnetic flux units in the system, and the other the effective strength of spin-orbit self-interaction. This latter interaction can be either attractive/focusing or repulsive/defocusing, and depends both on the intrinsic spin-orbit interaction and the relative length scale of the flux profile of the anyons. Densities and energies of ground and excited states are studied analytically and numerically for a wide range of the parameters and align well with a sequence of exact nonlinear Landau levels describing Jackiw–Pi self-dual solitons. With increasing flux, counter-rotating vortices are formed, enhancing the stability of the gas against collapse. Apart from clarifying the relations between various different anyon models that have appeared in the literature, our analysis sheds new light on the many-anyon spectral problem, and also exemplifies a novel supersymmetry-breaking phenomenon.
@@
$\int_{D(0,1)}\left[|\nabla f|^{2}+\alpha^{2}\frac{|\mathbf{x}|^{2}}{\max\{R,|\mathbf{x}|\}^{4}}|f|^{2}+\alpha g\frac{{\mathds{1}}_{D(0,R)}}{R^{2}}|f|^{2}\right],$
@@
where $P,Q$ are linearly independent and coprime complex polynomials with $n:=\max(\deg P,\deg Q)=\beta/2$. This complete set of regular minimizers includes all the self-dual soliton solutions found by Jackiw and Pi: initially, radial vortex rings with $P=z^{n}$ and $Q=1$ JacPi-90b , and eventually, the more general rationals $P/Q$ JacPi-90a . Thus, along the critical line $\gamma=-2\pi\beta$, $\beta\in\mathbb{R}$, the sequence of points $\beta=2n$ marks a phase transition from strictly positive (broken supersymmetry) to zero energy and increasing degree of supersymmetry, exhibited by NLL manifolds of real dimension $4n$; compare Figs. 2 and 3 (top).
@@
For a general estimate of the ground-state energy at $\gamma=2\pi\beta\kappa$, $\kappa\geq-1$, we have compared with the linear interpolation $E^{\mathrm{TF}}(2\pi c(1+\kappa)\beta)=E^{\mathrm{TF}}(2\pi\beta)\sqrt{c}\sqrt{1+\kappa}$ (dash-dotted curve), which provides a fit to numerics to within a few percent for $\beta\gtrsim 3$ and $0.9\lesssim\kappa\lesssim 2$. The estimate given by ChiSen-92 for $\kappa=1$ is ca 7% smaller. An absolute lower bound for the energy (assuming regularity and singlevaluedness) is provided by $\gamma_{*}(\beta)\to 2\pi\beta$ and thus differs by a factor $\sqrt{c}$, i.e. 9% (dotted curve; which for ideal anyons is still 41% higher than the ideal Fermi energy). We also conjecture that the real-valued minimizer at $\kappa=\pm 1$ has a linearly interpolating energy $|2\pm\beta|$. This, as well as other linear features of the $\kappa=\pm 1$ spectra, resembles the known linear/noninterpolating exact states of the $N$-anyon spectrum Chou-91a ; MurLawBhaDat-92 ; ChiSen-92 ; AtaGirLun-25 .
@@ Acknowledgments.
We are grateful to Romain Duboscq for the original Matlab code upon which much of our numerical work is based. A.A., A.E., F.G. and D.L. thank the Department of Mathematics of Politecnico di Milano for its kind hospitality during the spring of 2025 intensive period “Quantum Mathematics at Polimi”. D.L. also thanks the GSSI L’Aquila for its hospitality. All of us have benefited from discussions at the Institut Mittag-Leffler workshop “Anyons from small to large scales” in July 2025. Financial support from the Swedish Research Council (D.L., grant no. 2021-05328, “Mathematics of anyons and intermediate quantum statistics”) is gratefully acknowledged. A.A. thanks Kaj Nyström for the financial support for traveling to the relevant conferences. The author T.G. acknowledges the Gran Sasso Science Institute for the financial support making this work possible. A.E. gratefully thanks the Center for Interdisciplinary Mathematics at Uppsala University as well as C.F. Liljewalchs stipendiestiftelse for their financial support. D.-T.N. was funded by the Postdoctoral Scholarship Programme of Vingroup Innovation Foundation (VINIF), VinUniversity, code VINIF.2025.STS.15, and also gratefully acknowledges the support of the Knut and Alice Wallenberg Foundation through the grant of Prof. Andreas Strömbergsson during his postdoctoral period at Uppsala University (2022–2024).
@@
Due to the determinantal property of the Wronskian and the symmetry of the amplitude $|P|^{2}+|Q|^{2}$, the corresponding NLL state $\phi_{P,Q}$ given by (8) is invariant under pointwise action of the group $\mathrm{SU}(2)\times\mathbb{R}^{+}$ on the space of polynomial pairs $(P,Q)$. It follows that we may without loss of generality assume that either $P$ or $Q$ is monic, and, e.g., that $\deg Q<\deg P=n=\beta/2$. We may also simplify the problem (15) by adding any multiple of $Q$ to $P$. If we identify states that differ only by multiplication by a constant phase factor $e^{i\theta}$, we have a further $\mathrm{U}(1)$-symmetry, for a full $\mathrm{U}(2)\times\mathbb{R}^{+}$-symmetry in total. See AtaLunNgu-24 for further details.
@@
with $w_{1},w_{2}\in\mathbb{C}$ arbitrary. If $w_{1}=w_{2}$, this is a vortex ring (19) of degree 2. For $d=3$, we have two particular solutions:
@@
For $d=4$, we have the two particular solutions
@@
The above particular solutions yield all NLL states for $n=3$ via (16), and $P,Q$ are coprime at least if all the roots $w_{j}$ are disjoint (consider $f^{\prime}=P^{\prime\prime}Q-PQ^{\prime\prime}$).
@@
First we make some general observations. By expanding the kinetic-magnetic term $\lvert(\nabla+i\beta\mathbf{A}_{\varrho})\phi\rvert^{2}$, the energy functional can be recast into the form
@@
which satisfies $\int_{\mathbb{R}^{2}}\varrho^{a}=1$. The corresponding magnetic potential is $\mathbf{A}_{\varrho^{a}}=A(r)\mathbf{e}_{\theta}$, with
@@
$\int_{\mathbb{R}^{2}}\lvert\mathbf{A}_{\varrho^{a}}\rvert^{2}\varrho^{a}=(2a-1)\vartheta(a),$
where we introduce
$\vartheta(a):=2\psi(4a-1)-\psi(2a)-\psi(6a-2).$ (46)
This can be shown by evaluating the integral
$\int_{\mathbb{R}^{2}}\lvert\mathbf{A}_{\varrho^{a}}\rvert^{2}\varrho^{a}=2(2a-1)\int_{0}^{\infty}\frac{(1-(1+r^{2})^{-(2a-1)})^{2}}{r(1+r^{2})^{2a}}\,dr,$
@@
$=\sqrt{\frac{2(2a-1)}{a-1}}\sqrt{\frac{2a}{2a+1}+\frac{1}{\pi}\frac{2a-1}{4a-1}\gamma+\vartheta(a)\beta^{2}}.$ (48)
For fixed $(\beta,\gamma)$, this expression can be numerically optimized over $a>1$, yielding an optimal state $\phi^{a_{*}}$. This has been done using Mathematica, and the resulting curve $E=\mathcal{E}[\phi^{a_{*}}]$ for $\gamma=2\pi\beta\kappa$ is the thick yellow line in the plots.
We note that the Gaussian can be re-obtained by e.g. setting $a=2/b$, $\lambda=\sqrt{b/2}$, and taking the limit $b\to 0$. Hence the state $\phi^{a}$ can (for optimal $a>1$) be expected to have a lower energy than the Gaussian state.
@@
Here we tabulate approximate energies of (2) computed in Matlab and optimized using a steepest descent algorithm adopted from CorDubLunRou-19 . As initial states we have started from the states generated in the work CorDubLunRou-19 with $\gamma=0$, and adjusted the parameters $(\beta,\gamma)$ to nearby values in order to generate approximate ground states with different vorticities (the number of vortices listed is approximate in non-symmetric cases). In the stable regime away from criticality, convergence to the desired precision is typically achieved after 50-200 iterations of the algorithm. Due to geometrically and topologically induced metastability, the energy converges into a local minimum, and the global minimum may only be found by comparing a set of reasonable vorticities. Some chosen initial states exhibit sufficient metastability that they can produce estimates for excited states well above the ground levels, by stopping the algorithm after a limited number of iterations during which a stable phase is shown. Thus, all numerically computed energies are only valid as approximate upper bounds to the true ground-state energy, but may also yield indications of higher resonances. At criticality $\kappa=-1$ we have only been able to obtain sufficiently stable and reliable numerics for radially symmetric states. We have also noted some observations, such as approximate linearity, comparison to analytical energies [^1] Note that the constant $c=2\sqrt{\pi}/3$ is a numerical fit approximation to that denoted by $e(1,1)/(2\pi)$ in the earlier work CorDubLunRou-19 ., or cases where we suspect that the state is far from converged. The resulting energies are shown in the plots using solid markers (colors and shapes indicate numbers of visible vortices; see legend in Fig. 4 (bottom)).